OpenStudy (anonymous):

Translate the sentence into an inequality. Eight times the sum of a number and 27 is at least 25. Use the variable x for the unknown number.

OpenStudy (lieutenantgeneral):

THIS IS EASY

OpenStudy (lieutenantgeneral):

(n + 27) * 8

OpenStudy (lieutenantgeneral):

27 * a fraction = 25 let me do it

OpenStudy (lieutenantgeneral):

27 * 25/27 = 25

OpenStudy (anonymous):

\[8 (n+27)\leq 25 \]

OpenStudy (lieutenantgeneral):

oh ur doing inequalities? OHH WHY DIDN'T THE PERSON TELL ME